I am having trouble setting up the equation to solve this problem. A bag contains 42 coins, with a total weight of 246 grams. if the bag contains only gold coins that weigh 8 grams each and silver coins that weigh 5 grams each, how many gold and how many silver coins are in the bag?

Answer by htmentor(1343)

Let g = the number of gold coins weighing 8 g
Then 42 - g = the number of silver coins weighing 5 g
We can write the following expression for the total weight:
246 = 8g + 5(42-g)
Simplify and solve for g:
246 = 8g + 210 - 5g
3g = 36
g = 12
So there are 12 gold coins and 30 silver coins
Check:
12*8 + 30*5 = 96 + 150 = 246